The Pencil Code Newsletter

Issue 2023/3

October 21, 2023, Revision: 1.6

Contents

1 Steering committee election  1
2 Pencil Code User Meeting  1
3 Proposals for PCUM in 2025  2
4 Reports from the last PCUM  2
5 Code developments  2
5.1 Saving memory for df . . . . . . . . . . 2
6 Meetings  2
6.1 Early career meeting in Salzburg …. 2
7 Papers since April 2023  2

1 Steering committee election

We have currently 20 owners; most of them have declared themselves being available for serving on the Pencil Code Steering Committee (PCSC);

see http://norlx65.nordita.org/~brandenb/penci

l-code/PCSC/ for the terms of reference and the minutes of last meetings.

In the following, we list those eligibe to be voted into the next PCSC in alphabetic order. Many of them have provided a paragraph about themselves.

Philippe Bourdin. My start with the Pencil Code (PC) was 15 years ago and I gained experience in solar physics and also on the technical side of the code due to my demands in high-performance computing. Since then, I introduced several mechanisms to let the PC scale on more CPUs. In the past 2.5 years, I served as a PCSC member and was elected as chair of the PCSC some months ago. I am very grateful for this proof of trust and would like to candidate again as PCSC member. My aim is to incorporate my technical understanding in the decision making process of the PC community, continue supporting the community as a whole, as well as being a contact person for PC users and owners. Because the task of the PCSC is to deal with possible personal conflicts, I like to emphasize that I address such conflicts calmly and professionally. I look forward to supporting the community with my experience and patience. Thank you.

Axel Brandenburg. Together with Wolfgang Dobler, we started the Pencil Code as a community project from the very beginning, before anything worked – or even compiled. Now, 22 years later, it is a pleasure to see how the code has grown to have an h-index of 37, with 37 users having done at least

37 commits (https://github.com/pencil-code/p
encil-code/graphs/contributors). With now 20

owners, who can give check-in permissions to others, the code can stand on its owb – even when the original roots get old. One of my main ambitions is to facilitate maximum freedom in the sense that everything is allowed – as long as it doesn’t disturb anybody else. Of course, the auto-test doesn’t capture everything, and there are also a number of rules to prevent things going out of hand. This is why we have the PCSC, so nobody is allowed to take critical decisions alone, and bigger things are decided upon on behalf of the community. The PCSC reports back to the voters and users. For the next three years, if I’m elected, I continue to encourage users to make and check in their new developments for the benefit of the whole community.

Jennifer Schober. I am thrilled to put forth my candidacy as a committee member for the PCSC. My name is Jennifer Schober and I am a group leader at the Laboratoire d’Astrophysique, École Polytechnique Fédérale de Lausanne (EPFL), Switzerland. I started using the Pencil Code during my first postdoctoral fellowship at Nordita in 2015 and became a code owner in 2022. My main interest is MHD turbulence in different astrophysical systems, from the early Universe, to galaxy clusters, and galaxies. In particular, I use the Pencil Code to explore the characteristics of MHD dynamos, delving into their nonlinear properties and uncovering essential details. If elected, I am eager to collaborate with fellow committee members to ensure the Pencil Code remains robust, adaptable, and in harmony with the evolving needs of our community.

2 Pencil Code User Meeting

September 23, Monday. Therefore, we will start planning to have the next Pencil Code User Meeting from September 23 through September 27 at the Institute of Space Sciences in Bellaterra (Barcelona).

  Barcelona Doodle on: https://doodle.com/meeti

ng/participate/id/dG8yMN7e

pictures?

3 Proposals for PCUM in 2025

For future ...

4 Reports from the last PCUM

lmodify Since 19 years in the manual.

5 Code developments

As usual, we find the latest code developments

on https://github.com/pencil-code/pencil-code

/commits/master. The green and red check marks indicate whether the travis test went through or not;

see https://app.travis-ci.com/github/pencil-co
de/pencil-code.

5.1 Saving memory for df

6 Meetings

6.1 Early career meeting in Salzburg

Philippe ....

the Austrian Society for Astronomy and Astrophysics (GAA) organizes a conference mainly aimed for ”early career scientists” with the focus to foster networks among Master students, PhD students, and young PostDocs. Supervisors are invited to participate, too.

We like to give you the opportunity to apply for a talk in a real conference environment, while you are still at an early career stage. Colleagues from Munich, Vienna, Innsbruck, and Graz will participate. As the venue is in Salzburg/Austria, it is easily reachable in Europe.

The topics include (but are not limited to) astronomy, astrophysics, astroparticle physics, cosmology, solar and heliospheric physics, space sciences, as well as exoplanets, planets, and small bodies.

Date: 8th and 9th of March, 2024 (noon to noon) Registration: now open, until 1st of December Fees: 10 (for non-GAA members) More details and registration: https://www.oegaa.at/AECC/ (Transportation to the venue from Salzburg city will be organized.)

Thank you for your kind consideration, on behalf of the GAA board, Philippe Bourdin.

7 Papers since April 2023

As usual, we look here at new papers that make use of the Pencil Code. Since the last newsletter of April 1st, 14 new papers have appeared on the arXiv, and 11 others, some of which were just preprints and have now been published. We list both here, 25 altogether. A browsable ADS list of all Pencil Code papers

                                               can be found on: https://ui.adsabs.harvard.ed
                                               u/public-libraries/iGR7N570Sy6AlhDMQRTe_A. If

something is missing in those entries, you can also

                                               include it yourself in: https://github.com/pencil-

                                               code/pencil-code/blob/master/doc/citations/re

f.bib, or otherwise just email brandenb@nordita.org. A compiled version of this file is available as https:/

                                               /github.com/pencil-code/website/blob/master/d

oc/citations.pdf, where we also list a total of now 102 code comparison papers in the last section “Code comparison & reference”. Those are not included in our list below, nor among the now total number of 630 research papers that use the Pencil Code.

References

Brandenburg, A., Clarke, E., Kahniashvili, T., Long, A.J. and Sun, G., Relic Gravitational Waves from the Chiral Plasma Instability in the Standard Cosmological Model. arXiv e-prints, 2023a, arXiv:2307.09385.

Brandenburg, A., Kamada, K., Mukaida, K., Schmitz, K. and Schober, J., Chiral magnetohydrodynamics with zero total chirality. Phys. Rev. D, in press, 2023b, arXiv:2304.06612.

Brandenburg, A., Kamada, K. and Schober, J., Decay law of magnetic turbulence with helicity balanced by chiral fermions. Phys. Rev. Res., 2023c, 5, L022028.

Brandenburg, A. and Larsson, G., Turbulence with Magnetic Helicity That Is Absent on Average. Atmosphere, 2023, 14, 932.

Brandenburg, A. and Protiti, N.N., Electromagnetic conversion into kinetic and thermal energies. Entropy, 2023, 25, 1270.

Brandenburg, A., Sharma, R. and Vachaspati, T., Inverse cascading for initial MHD turbulence spectra between Saffman and Batchelor. arXiv e-prints, 2023d, arXiv:2307.04602.

Candelaresi, S. and Beck, C., Twisted magnetic knots and links. Phys. Plasmas, 2023, 30, 082102.

Elias-López, A., Del Sordo, F. and Viganò, D., Vorticity and magnetic dynamo from subsonic expansion waves. arXiv e-prints, 2023, arXiv:2304.11929.

Ganti, H., Bravo, L. and Khare, P., Interactions between high hydrogen content syngas-air premixed flames and homogeneous isotropic turbulence: Flame thickening. Phys. Fluids, 2023, 35, 075150.

Gent, F.A., Mac Low, M.M. and Korpi-Lagg, M.J., Transition from small-scale to large-scale dynamo in a supernova-driven, multiphase medium. arXiv eprints, 2023, arXiv:2306.07051.

Hackman, T., Kochukhov, O., Viviani, M., Warnecke, J., Korpi-Lagg, M.J. and Lehtinen, J.J., From convective stellar dynamo simulations to Zeeman-Doppler images. arXiv e-prints, 2023, arXiv:2306.07838.

He, Y., Roper Pol, A. and Brandenburg, A., Modified propagation of gravitational waves from the early radiation era. J.. Cosmol. Astropart. Phys., 2023, 2023, 025.

Käpylä, P.J., Browning, M.K., Brun, A.S., Guerrero, G. and Warnecke, J., Simulations of solar and stellar dynamos and their theoretical interpretation. arXiv e-prints, 2023, arXiv:2305.16790.

Karak, B.B., Models for the long-term variations of solar activity. Liv. Rev. Solar Phys., 2023, 20, 3.

Lipatnikov, A.N. and Sabelnikov, V.A., Influence of small-scale turbulence on internal flamelet structure. Phys. Fluids, 2023, 35, 055128.

Lyra, W., Stability Analysis for General Order Central Finite-difference Hyperdiffusivity with Time Integrators of Arbitrary Accuracy. Res. Not. Am. Astron. Soc., 2023, 7, 69.

Mizerski, K.A., Yokoi, N. and Brandenburg, A., Crosshelicity effect on α-type dynamo in non-equilibrium turbulence. J. Plasma Phys., 2023, 89, 905890412.

Mondal, T. and Bhat, P., A unified treatment of mean-field dynamo and angular-momentum transport in magnetorotational instability-driven turbulence. arXiv e-prints, 2023, arXiv:2307.01281.

Navarrete, F.H., Käpylä, P.J., Schleicher, D.R.G. and Banerjee, R., Effects of the centrifugal force on stellar dynamo simulations. arXiv e-prints, 2023, arXiv:2305.01312.

Ortiz-Rodríguez, C.A., Käpylä, P.J., Navarrete, F.H., Schleicher, D.R.G., Mennickent, R.E., Hidalgo, J.P. and Toro, B., Simulations of dynamo action in slowly rotating M dwarfs: Dependence on dimensionless parameters. arXiv e-prints, 2023, arXiv:2305.16447.

Pavaskar, P., Yan, H. and Cho, J., Magnetic field measurement from the Davis-Chandrasekhar-Fermi method employed with atomic alignment. Month. Not. Roy. Astron. Soc., 2023, 523, 1056–1066.

Schober, J., Rogachevskii, I. and Brandenburg, A., Chiral magnetic anomaly and dynamos from spatial chemical potential fluctuations. arXiv e-prints, 2023, arXiv:2307.15118.

Tharakkal, D., Shukurov, A., Gent, F.A., Sarson, G.R. and Snodin, A., Steady states of the Parker instability: the effects of rotation. arXiv e-prints, 2023, arXiv:2305.03318.

Warnecke, J., Korpi-Lagg, M.J., Gent, F.A. and Rheinhardt, M., Numerical evidence for a small-scale dynamo approaching solar magnetic Prandtl numbers. Nat. Astron., 2023.

Zhu, J.Z. and Shi, P.X., Helical and nonhelical (magneto-)Burgers turbulence: I. Compressibility reduction and beyond. arXiv e-prints, 2023, arXiv:2307.05490.

This Pencil Code Newsletter was edited by

                                             Axel  Brandenburg  <brandenb@nordita.org>,

Nordita, KTH Royal Institute of Technology and Stockholm University, SE-10691 Stockholm, Sweden; and Matthias Rheinhardt

                                             <matthias.rheinhardt@aalto.fi>,  Department

of Computer Science, Aalto University, PO Box 15400, FI-00076 Aalto, Finland. See http://www.nordita.

                                             org/~brandenb/pencil-code/newsletter or https

                                             ://github.com/pencil-code/website/tree/master

/NewsLetters for the online version as well as back issues.